If I invest 11000 dollars for 5 years at 11% compounded quarterly, how much would i have?
Answer by Nate(3500)

P(a) = a(1 + r/n)^(tn)
P(a) = 11000(1 + .11/4)^20
P(a) = 11000(1.0275)^20
P(a) = 18924.71